# DeepLight: Reconstructing High-Resolution Observations of Nighttime Light With Multi-Modal Remote Sensing Data

### v2

@@ Abstract
Nighttime light (NTL) remote sensing observation serves as a unique proxy for quantitatively assessing progress toward meeting a series of Sustainable Development Goals (SDGs), such as poverty estimation, urban sustainable development, and carbon emission. However, existing NTL observations often suffer from pervasive degradation and inconsistency, limiting their utility for computing the indicators defined by the SDGs. In this study, we propose a novel approach to reconstruct high-resolution NTL images using multi-modal remote sensing data. To support this research endeavor, we introduce DeepLightMD, a comprehensive dataset comprising data from five heterogeneous sensors, offering fine spatial resolution and rich spectral information at a national scale. Additionally, we present DeepLightSR, a calibration-aware method for building bridges between spatially heterogeneous modality data in the multi-modality super-resolution. DeepLightSR integrates calibration-aware alignment, an auxiliary-to-main multi-modality fusion, and an auxiliary-embedded refinement to effectively address spatial heterogeneity, fuse diversely representative features, and enhance performance in 8×8$\times$8 × super-resolution (SR) tasks. Extensive experiments demonstrate the superiority of DeepLightSR over 8 competing methods, as evidenced by improvements in PSNR (2.01 dB $\sim$ 13.25 dB) and PIQE (0.49 $\sim$ 9.32). Our findings underscore the practical significance of our proposed dataset and model in reconstructing high-resolution NTL data, supporting efficiently and quantitatively assessing the SDG progress.
@@ 1 Introduction
The pursuit of the Sustainable Development Goals (SDGs) UnitedNations (2015) encounters various hurdles such as data insufficiency and lack of quantitative progress assessment Xi et al. (2023). Earth observation data from satellites emerges as a promising solution by providing large-scale quantitative attributes and capturing detailed spatial information essential for understanding Earth’s ecosystem, thus being recognized as a highly relevant tool in accomplishing SDGs Guo et al. (2023).
Observations of nighttime light (NTL) over several decades effectively characterize both the extent and intensity of human activities, providing unique insight for quantitatively assessing the progress toward the SDGs Maso et al. (2023), such as investigating poverty and inequality McCallum et al. (2022), monitoring urban sustainable developments Elvidge et al. (1997), and estimating carbon emissions Huang et al. (2014); Levin et al. (2020). Due to the lack of onboard calibration, sensor degradation, and satellite orbit drift, the existing historical record of low-resolution NTL suffers from severe quality degradation like blurring, blooming effect, and over-saturation phenomenon Abrahams et al. (2018); Hsu et al. (2015); Li et al. (2020b), highlighting the urgent need for improving the historical NTL data collection. To alleviate these problems, a cost-effective way is to introduce super-resolution (SR) techniques, which reconstruct high-resolution (HR) NTL images from low-resolution (LR) NTL images.
However, as illustrated in Fig. 1, the degradations in historical LR NTL observations are composed of spatially heterogeneous inconsistency, coarse resolution, sensor blurring, atmospheric scattering, etc. The complexity of degradation types makes it a crucial task to improve the quality of LR NTL data for over two decades Falchi et al. (2016). Therefore, We present the task of reconstructing HR NTL with multiple modalities in this study. Given an LR NTL image, the algorithm is expected to fuse multi-modality features to guide the reconstruction of calibrated HR NTL image. Nevertheless, the LR images in existing SR datasets are usually manual-simulated with common degradations such as Gaussian blur Dong et al. (2022). Considering the generalization of the model in addressing various degradations relies on the representativeness of SR datasets, the existing SR datasets are inadequate to support the training for this task.
@@ 2.1 Calibration-Based Methods in Remote Sensing
Assuming that there is a place where ground truth NTL intensity seldom changed over the past several decades, the main theory of calibration-based methods is to calibrate any pairs of cross-sensor NTL images from one to another using this invariant target. Lots of efforts have been paid to reconstructing calibrated NTL dataset with a single invariant target Elvidge et al. (2009), several invariant targets Wu et al. (2013), and invariant regions Wang et al. (2022). Apart from expert knowledge, another way relies on determining invariant pixels using machine learning and deep-learning algorithms. For example, A ridge sampling and regression method is proposed to calibrate the NTL time series over the entire globe Zhang et al. (2016). An embedded convolutional neural network (CNN) framework is proposed to calibrate NTL images Chen et al. (2021). Modified U-Nets were introduced to achieve NTL images from different remote sensing modalities Zhang et al. (2021b); Nechaev et al. (2021). However, the existing calibration-based methods are proposed to calibrate NTL images with ×$\times 2$ scaling factor and find it challenging to address most degradation issues in DMSP-OLS NTL data, resulting in limitations of practical and extended applications.
@@ 2.2 Super-Resolution -Based Methods in Computer Vision
Quantities of deep learning methods have been proposed in effectively conducting super-resolution (SR) tasks in CV, such as SRCNN Dong et al. (2016), SRResNet Ledig et al. (2017), and RCAN Haut et al. (2019). However, most existing SR methods are designed based on the assumption that LR input and HR target lie in the same domain, which is not valid in this task. We summarize cross-modality SR methods which are or can potentially be applied in applications where LR and HR belong to different domains. The first category is the single-modality SR methods, which do not involve auxiliary modality in the reconstruction of SR results, such as unpaired LR-HR under different domains Maeda (2020). Besides, several efforts have been paid to different SR targets, such as magnetic resonance (MR) images Fang et al. (2022), depth images Sun et al. (2021), and thermal images Gupta and Mitra (2022). In order to improve the reconstruction performance, auxiliary modalities are introduced to provide useful information to guide the reconstruction. For example, HR RGB visual modality is introduced in the reconstruction of non-visual data Shacht et al. (2021). JMDL model is proposed to solve the general SR problem with one auxiliary modality Deng and Dragotti (2020). However, the existing multi-modality SR methods are designed to alleviate simple degradation and weak misalignment problem in conducting SR tasks, which are difficult to address the complicated degradation and spatially heterogeneous inconsistency of multi-modality NTL SR tasks.
@@ 3.1 Data Collection
HR NTL: The composed Luojia NTL data Li et al. (2019) is selected as HR NTL in DeepLightMD. The Luojia-01 provides global NTL images at 130 m with 15 days as revisit time. Compared to other commonly used NTL datasets such as VIIRS, the Luojia-01 provides NTL observations with higher spatial resolution, more detailed information, and a wider radiometric range. Besides, the overpass time gap between DMSP and Luojia-01 is much smaller than any other publicly available NTL pairs, reducing the challenges posed by overpass time misalignment Li et al. (2020a).
@@ 6 Conclusion and Future Work
In this paper, we introduce DeepLightMD, a pioneering national-scale benchmark dataset encompassing data from five modalities. Building upon this dataset, we propose DeepLightSR, a calibration-aware approach designed to reconstruct HR NTL data. Both qualitative and quantitive results on DeepLightMD demonstrate the effectiveness of the proposed DeepLightSR method in generating large-scale, high-resolution NTL data. In the future, Our work holds promise in providing more precise NTL data which can be instrumental in mapping urbanization processes and monitoring street-level greenhouse gas emissions. By contributing to SDG targets related to poverty alleviation, sustainable urban development, and climate action, our research paves the way for positive societal impact and progress assessment towards the SDGs.
Furthermore, we identify several avenues for future research with potential societal impact aligned with SDGs: 1) Consistent long-term NTL data reconstruction. There is a growing demand to extend our methodology to reconstruct consistent long-term NTL datasets, addressing temporal variability while ensuring alignment. 2) Incorporating more modalities in the reconstruction of NTL images. Incorporating additional modalities, such as Interferometric Synthetic Aperture Radar (SAR) and social media data, holds promise in enhancing performance by capturing more nuanced aspects related to terrain and socioeconomic factors.

### v3

@@ Abstract
Nighttime light (NTL) remote sensing observation serves as a unique proxy for quantitatively assessing progress toward meeting a series of Sustainable Development Goals (SDGs), such as poverty estimation, urban sustainable development, and carbon emission. However, existing NTL observations often suffer from pervasive degradation and inconsistency, limiting their utility for computing the indicators defined by the SDGs. In this study, we propose a novel approach to reconstruct high-resolution NTL images using multi-modal remote sensing data. To support this research endeavor, we introduce DeepLightMD, a comprehensive dataset comprising data from five heterogeneous sensors, offering fine spatial resolution and rich spectral information at a national scale. Additionally, we present DeepLightSR, a calibration-aware method for building bridges between spatially heterogeneous modality data in the multi-modality super-resolution. DeepLightSR integrates calibration-aware alignment, an auxiliary-to-main multi-modality fusion, and an auxiliary-embedded refinement to effectively address spatial heterogeneity, fuse diversely representative features, and enhance performance in 8×8$\times$8 × super-resolution (SR) tasks. Extensive experiments demonstrate the superiority of DeepLightSR over 8 competing methods, as evidenced by improvements in PSNR (2.01 dB $\sim$ 13.25 dB) and PIQE (0.49 $\sim$ 9.32). Our findings underscore the practical significance of our proposed dataset and model in reconstructing high-resolution NTL data, supporting efficiently and quantitatively assessing the SDG progress. The code and data will be available at https://github.com/xian1234/DeepLight.
@@ 1 Introduction
The pursuit of the Sustainable Development Goals (SDGs) UnitedNations (2015) encounters various hurdles such as data insufficiency and lack of quantitative progress assessment Xi et al. (2023). Earth observation data from satellites emerges as a promising solution by providing large-scale quantitative attributes Zhang et al. (2024b) and capturing detailed spatial information essential for understanding Earth’s ecosystem, thus being recognized as a highly relevant tool in accomplishing SDGs Guo et al. (2023).
Observations of nighttime light (NTL) over several decades effectively characterize both the extent and intensity of human activities, providing unique insight for quantitatively assessing the progress toward the SDGs Maso et al. (2023), such as investigating poverty and inequality McCallum et al. (2022), monitoring urban sustainable developments Elvidge et al. (1997), and estimating carbon emissions Huang et al. (2014); Levin et al. (2020). Due to the lack of onboard calibration, sensor degradation, and satellite orbit drift, the existing historical record of low-resolution NTL suffers from severe quality degradation like blurring, blooming effect, and over-saturation phenomenon Abrahams et al. (2018), highlighting the urgent need for improving the historical NTL data collection. To alleviate these problems, a cost-effective way is to introduce super-resolution (SR) techniques, which reconstruct high-resolution (HR) NTL images from low-resolution (LR) NTL images.
However, as illustrated in Fig. 1, the degradations in historical LR NTL observations are composed of spatially heterogeneous inconsistency, coarse resolution, sensor blurring, atmospheric scattering, etc. The complexity of degradation types makes it a crucial task to improve the quality of LR NTL data for over two decades Falchi et al. (2016). Therefore, We present the task of reconstructing HR NTL with multiple modalities in this study. Given an LR NTL image, the algorithm is expected to fuse multi-modality features to guide the reconstruction of calibrated HR NTL image. Nevertheless, the LR images in existing SR datasets are usually manual-simulated with common degradations Dong et al. (2023). Considering the generalization of the model in addressing various degradations relies on the representativeness of SR datasets, the existing SR datasets are inadequate to support the training for this task.
@@ 2.1 Calibration-Based Methods in Remote Sensing
Assuming that there is a place where ground truth NTL intensity seldom changed over the past several decades, the main theory of calibration-based methods is to calibrate any pairs of cross-sensor NTL images from one to another using this invariant target. Lots of efforts have been paid to reconstructing calibrated NTL dataset with a single invariant target Elvidge et al. (2009), several invariant targets Wu et al. (2013), and invariant regions Wang et al. (2022). Apart from expert knowledge, another way relies on determining invariant pixels using machine learning and deep-learning algorithms. For example, A ridge sampling and regression method is proposed to calibrate the NTL time series over the entire globe Zhang et al. (2016). An embedded convolutional neural network (CNN) framework is proposed to calibrate NTL images Chen et al. (2021). Modified U-Nets were introduced to achieve NTL images from different remote sensing modalities Zhang et al. (2021b). However, the existing calibration-based methods are proposed to calibrate NTL images with ×$\times 2$ scaling factor and find it challenging to address most degradation issues in DMSP-OLS NTL data, resulting in limitations of practical and extended applications.
@@ 2.2 Super-Resolution -Based Methods in Computer Vision
Quantities of deep learning methods have been proposed in effectively conducting super-resolution (SR) tasks in CV Zhang et al. (2021a); Dong et al. (2024), such as SRCNN Dong et al. (2016), SRResNet Ledig et al. (2017), and RCAN Haut et al. (2019). However, most existing SR methods are designed based on the assumption that LR input and HR target lie in the same domain, which is not valid in this task. We summarize cross-modality SR methods which are or can potentially be applied in applications where LR and HR belong to different domains. The first category is the single-modality SR methods, which do not involve auxiliary modality in the reconstruction of SR results, such as unpaired LR-HR under different domains Maeda (2020). Besides, several efforts have been paid to different SR targets, such as magnetic resonance (MR) images Fang et al. (2022), depth images Sun et al. (2021), and thermal images Gupta and Mitra (2022). In order to improve the reconstruction performance, auxiliary modalities are introduced to provide useful information to guide the reconstruction. For example, HR RGB visual modality is introduced in the reconstruction of non-visual data Shacht et al. (2021). JMDL model is proposed to solve the general SR problem with one auxiliary modality Deng and Dragotti (2020). However, the existing multi-modality SR methods are designed to alleviate simple degradation and weak misalignment problem in conducting SR tasks, which are difficult to address the complicated degradation and spatially heterogeneous inconsistency of multi-modality NTL SR tasks.
@@ 3.1 Data Collection
HR NTL: The composed Luojia NTL data Li et al. (2019) is selected as HR NTL in DeepLightMD. The Luojia-01 provides global NTL images at 130 m with 15 days as revisit time. Compared to other commonly used NTL datasets such as VIIRS, the Luojia-01 provides NTL observations with higher spatial resolution, more detailed information, and a wider radiometric range. Besides, the overpass time gap between DMSP and Luojia-01 is much smaller than any other publicly available NTL pairs, reducing the challenges posed by overpass time misalignment Li et al. (2020).
@@ 6 Conclusion and Future Work
In this paper, we introduce DeepLightMD, a pioneering national-scale benchmark dataset encompassing data from five modalities. Building upon this dataset, we propose DeepLightSR, a calibration-aware approach designed to reconstruct HR NTL data. Both qualitative and quantitive results on DeepLightMD demonstrate the effectiveness of the proposed DeepLightSR in generating large-scale, high-resolution NTL data. Furthermore, we identify several avenues for future research with potential societal impact aligned with SDGs: 1) Consistent long-term NTL data reconstruction. There is a growing demand to extend our methodology to reconstruct consistent long-term NTL datasets, addressing temporal variability while ensuring alignment. 2) Incorporating more modalities in the reconstruction of NTL images. Incorporating additional modalities, such as Synthetic Aperture Radar (SAR) and social media data, holds promise in enhancing performance by capturing more nuanced aspects related to terrain and socioeconomic factors.
Acknowledgements
This research was supported in part by the National Natural Science Foundation of China (Grant No. T2125006 and No. 42301390), Jiangsu Innovation Capacity Building Program (Project No. BM2022028), China Postdoctoral Science Foundation (Grant No. 2023M731871).